OpenStudy (anonymous):

How do you solve -35=-15p

OpenStudy (anonymous):

You try to get p alone on one side

OpenStudy (dls):

Divide by -5 on both sides.

OpenStudy (shamil98):

-15 p = -35 you want to isolate p. so you divide -15 to both sides p = -35/-15 then reduce that fraction. p = 35/15 = 7/3